# Training-Free Safe Denoisers for Safe Use of Diffusion Models

### v2

@@ Abstract
There is growing concern over the safety of powerful diffusion models (DMs), as they are often misused to produce inappropriate, not-safe-for-work (NSFW) content or generate copyrighted material or data of individuals who wish to be forgotten. Many existing methods tackle these issues by heavily relying on text-based negative prompts or extensively retraining DMs to eliminate certain features or samples. In this paper, we take a radically different approach, directly modifying the sampling trajectory by leveraging a negation set (e.g., unsafe images, copyrighted data, or datapoints needed to be excluded) to avoid specific regions of data distribution, without needing to retrain or fine-tune DMs. We formally derive the relationship between the expected denoised samples that are safe and those that are not safe, leading to our safe denoiser which ensures its final samples are away from the area to be negated. Inspired by the derivation, we develop a practical algorithm that successfully produces high-quality samples while avoiding negation areas of the data distribution in text-conditional, class-conditional, and unconditional image generation scenarios. These results hint at the great potential of our training-free safe denoiser for using DMs more safely.
Machine Learning, ICML
@@ 1 Introduction
Diffusion models (DMs) have emerged as a powerful class of generative models, consistently surpassing previous approaches on a variety of tasks, including text-to-image generation (Rombach et al., 2022), audio synthesis (Kong et al., 2021), video synthesis (Bar-tal et al., 2024), and protein design (Watson et al., 2023). A significant factor contributing to their success is the flexible and controllable sampling with guidance (Dhariwal & Nichol, 2021; Ho & Salimans, 2021). In particular, text-based guidance (Saharia et al., 2022) has played a key role in the success of modern text-to-image models (Rombach et al., 2022; Podell et al., 2024).
Despite remarkable advancements, there is growing concern about the safety of content generated by these models. The first concern is regarding not-safe-for-work (NSFW) content generation. To tackle the concern, negative prompts (Gandikota et al., 2023a; Ban et al., 2024) have predominantly been used to guide models away from toxic text descriptions. Fine-tuning methods aimed at unlearning undesirable features (Gandikota et al., 2023a, b; Gong et al., 2024; Kim et al., 2024) have shown promise, too. However, their effectiveness is limited by adversarial attacks or jailbreaks that can circumvent safeguards (Zhang et al., 2024b; Yang et al., 2024).
Other safety concerns include DMs’ generation of copyrighted content and data of individuals who wish to be excluded (machine unlearning). These two concerns are closely related to DMs’s exceptional ability to memorize training data (Carlini et al., 2023). While differentially private training (Dockhorn et al., 2023; Liu et al., 2024) could mitigate the danger of memorization, there is an inevitable performance drop due to the added noise to the training process.
(a) Nudity (Top) &\&& Violence (Bottom)
(b) Data Memorization
Figure 1: Our method Safe Denoiser against existing methods. (a) Our method, incorporated with SLD (Schramowski et al., 2023) and SAFREE (Yoon et al., 2024), does not generate inappropriate images. (b) Our method mitigates the memorization issue by negating the real image, resulting in a novel image that contains similar features like grey hair and formal outfit to those in the real image.
In this work, we propose directly modifying the sampling trajectories of DMs such that the sampling trajectories adhere to theoretically safe distributions. The modification follows, what-we-call, safe denoiser, which is derived from the relationship (in Theorem. 3.2) between the expected denoised samples that are safe and those that are not safe. The final samples from the safe denoiser are theoretically guaranteed to be safe and away from the area to be negated. Based on this derivation, we develop a practical algorithm (in Algorithm 1) that approximates the theoretically safe denoiser to generate safe images or combined with existing negative prompting methods to enhance the safety of text-to-image generation.
In our experiments, we demonstrate that our safe denoiser achieves state-of-the-art performance in terms of its safe generation, in the tasks of concept erasing (a popular benchmark for avoiding NSFW images in text-to-image generation), class removal (object unlearning, a form of machine unlearning, in class conditional generation), and unconditional image generation. In what comes next, we start by describing background information.
2 Preliminary
@@
DMs generate samples through iterative decoding starting from random noise to data. This iterative process is a reverse of the forward data destruction process, given by $\mathbf{x}_{t}=\alpha_{t}\mathbf{x}+\sigma_{t}\bm{\epsilon}$, where $\mathbf{x}$ follows the data distribution $p_{\text{data}}(\mathbf{x})$ and $\bm{\epsilon}$ follows the noise prior distribution $\mathcal{N}(0,I)$, which results in the perturbation kernel to be $q_{t}(\mathbf{x}_{t}|\mathbf{x})=\mathcal{N}(\mathbf{x}_{t};\alpha_{t}\mathbf{x},\sigma_{t}^{2}I)$. The specific choice of the coefficients $\alpha_{t}$ and $\sigma_{t}$ determine the variant of DMs. Depending on these parameters, the model may be referred to as Denoising Diffusion Probabilistic Models (DDPM) (Ho et al., 2020), Elucidating Diffusion Models (EDM) (Karras et al., 2022), or flow matching (Lipman et al., 2022).
Regardless of whether the model is trained with noise-prediction (Ho et al., 2020), data-prediction (Karras et al., 2022), or velocity-prediction (Salimans & Ho, 2022; Lipman et al., 2022), these approaches are fundamentally equivalent (Kingma et al., 2021; Kim et al., 2021). This paper adopts the data-prediction framework due to its most intuitive interpretation. In data-prediction, the model approximates the denoiser function, defined by
$\begin{split}&\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]:=\int\mathbf{x}\frac{p_{\text{data}}(\mathbf{x})q_{t}(\mathbf{x}_{t}|\mathbf{x})}{p_{\text{data},t}(\mathbf{x}_{t})}\mathop{}\!\mathrm{d}\mathbf{x}\\ &\quad\quad\approx\frac{1}{\alpha_{t}}\big{(}\mathbf{x}_{t}+\sigma_{t}^{2}\mathbf{s}_{\bm{\theta}}\big{)}=\frac{1}{\alpha_{t}}\big{(}\mathbf{x}_{t}-\sigma_{t}\bm{\epsilon}_{\bm{\theta}}\big{)},\end{split}$ (1)
where $p_{\text{data},t}(\mathbf{x}_{t})$ is a marginal distribution of the noisy data distribution at time $t$, and $\mathbf{s}_{\bm{\theta}}$ and $\bm{\epsilon}_{\bm{\theta}}$ are score-prediction and noise-prediction, respectively.
@@
DMs can be guided to produce samples (Dhariwal & Nichol, 2021; Kim et al., 2022) that adhere more closely to a desired condition denoted by $\mathbf{c}$. A common approach in modern DMs is classifier-free guidance (CFG) (Ho & Salimans, 2021). The model is trained to learn both the unconditional denoiser $\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]$ and the conitional denoiser $\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t},\mathbf{c}]$. The CFG modifies the sampling trajectory by
$\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]+\lambda\big{(}\underbrace{\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t},\mathbf{c}]}_{\text{positive}}-\underbrace{\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]}_{\text{uncond}}\big{)}$
@@
allowing stronger alignment of the sample with the prompt $\mathbf{c}$ via the scale $\lambda$. The purpose of the additional term is to guide the unconditioinal denoiser in the sharpening direction toward a desired condition $\mathbf{c}$.
@@
Negative prompting (Liu et al., 2022) reverses the CFG gradient direction for an undesired prompt denoted by $\mathbf{c}_{-}$. Formally, one replaces the standard CFG update with
@@
where $\mathbf{c}_{+}$ denotes a positive condition and $\mathbf{c}_{-}$ represents a negative context, such as low quality, watermark, logo, etc.
Recently, Schramowski et al. (2023) introduced Safe Latent Diffusion (SLD), a new type of guidance, given by
$\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]+\underbrace{\lambda(\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t},\mathbf{c}_{+}]-\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]}_{\text{CFG}})$ (2)
$-\underbrace{\mu(\mathbf{c}_{+},\mathbf{c}_{US};\gamma,\lambda)(\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t},\mathbf{c}_{US}]-\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]}_{\text{SLD}}),$
where $\mathbf{c}_{US}$ represents a predefined set of unsafe prompts suggested by authors. Hypothetically, if we assume $\mu$ was set to be $\lambda$, the SLD guidance simplifies to
$\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]+\lambda(\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t},\mathbf{c}_{+}]-\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t},\mathbf{c}_{US}]).$
Instead of directly using $\mathbf{c}_{US}$ as $\mathbf{c}_{-}$, SLD introduces an adaptive weight $\mu(\mathbf{c}_{+},\mathbf{c}_{US};\gamma,\lambda)$ proportional to the denoiser difference norm, defined as $\|\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t},\mathbf{c}_{+}]-\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t},\mathbf{c}_{US}]\|$. The magnitude of this norm serves as an indicator of the proximity of the sampling trajectory to the unsafe region. Specifically, a larger norm suggests that the trajectory is likely to be safe, whereas a smaller norm indicates potential unsafety.
3 Methodology
The negative prompt $\mathbf{c}_{-}$ or the SLD prompt $\mathbf{c}_{US}$ consist of a limited set of pre-selected words by humans, and therefore may not encompass all images intended to be negated. Consequently, instead of ensuring safety solely based on text prompt, we introduce a methodology that guarantees safety based on images, which operates orthogonally to existing text-based safety approaches. Furthermore, while text-based negative guidance can enhance safety, its application lacks a theoretical foundation, thereby offering no guarantees regarding the distribution of the samples. To address these issues, we propose constructing a sampling trajectory that adheres to the safe distribution by using a safe denoiser defined below.
3.1 Safe Denoiser
@@
Figure 3: Effect of the weight value in Theorem 3.2. (a) If we use half the theoretical weight value, samples generated by our weak safe denoiser also cover the unsafe region (i.e., red dots appearing in the blue area). (b) When we use the theoretical value, the samples avoid unsafe regions while covering the whole safe area. (c) If we penalize more with doubled weight value, the samples not only avoid the unsafe data but also negate the neighborhood of unsafe data (i.e., there are no red dots in the black area).
@@
To define the safe denoiser, we first define an indicator function, $1_{\text{safe}}(\mathbf{x})$ taking the value of $1$ if $\mathbf{x}$ is safe and 00 if not. Similarly, we define an indicator function, $1_{\text{unsafe}}(\mathbf{x})$ taking the value of $1$ if $\mathbf{x}$ is unsafe and 00 if not. Hence, for each sample $\mathbf{x}$, we have a constant function, taking the value of $1$, defined by $1(\mathbf{x})=1_{\text{safe}}(\mathbf{x})+1_{\text{unsafe}}(\mathbf{x})$. Then, we define the following concepts.
Definition 3.1.
The unnormalized safe distribution $p_{\text{safe}}(\mathbf{x})$ is $1_{\text{safe}}(\mathbf{x})p_{\text{data}}(\mathbf{x})$. The safe denoiser is defined by
$\mathbb{E}_{\text{safe}}[\mathbf{x}|\mathbf{x}_{t}]=\int\mathbf{x}\frac{p_{\text{safe}}(\mathbf{x})q_{t}(\mathbf{x}_{t}|\mathbf{x})}{p_{\text{safe},t}(\mathbf{x}_{t})}\mathop{}\!\mathrm{d}\mathbf{x},$
@@
where $p_{\text{safe},t}(\mathbf{x}_{t})$ is the marginal distribution of the noisy safe data at $t$. Analogously, the unnormalized unsafe distribution $p_{\text{unsafe}}(\mathbf{x})$ is $1_{\text{unsafe}}(\mathbf{x})p_{\text{data}}(\mathbf{x})$. The unsafe denoiser is
$\mathbb{E}_{\text{unsafe}}[\mathbf{x}|\mathbf{x}_{t}]=\int\mathbf{x}\frac{p_{\text{unsafe}}(\mathbf{x})q_{t}(\mathbf{x}_{t}|\mathbf{x})}{p_{\text{unsafe},t}(\mathbf{x}_{t})}\mathop{}\!\mathrm{d}\mathbf{x},$ (3)
where $p_{\text{unsafe},t}(\mathbf{x}_{t})$ is the marginal distribution of the noisy unsafe data at $t$.
@@
Our interest is to obtain $\mathbb{E}_{\text{safe}}[\mathbf{x}|\mathbf{x}_{t}]$ given the data denoiser $\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]$ defined in Eq. (1). The theorem below describes the relationship between our safe denoiser and the usual data denoiser.
Theorem 3.2.
$\mathbb{E}_{\textup{safe}}[\mathbf{x}|\mathbf{x}_{t}]$ $=\mathbb{E}_{\textup{data}}[\mathbf{x}|\mathbf{x}_{t}]$ (4)
for a weight is defined by
$\beta^{*}(\mathbf{x}_{t})=\frac{Z_{\textup{unsafe}}p_{\textup{unsafe},t}(\mathbf{x}_{t})}{Z_{\textup{safe}}p_{\textup{safe},t}(\mathbf{x}_{t})},$ (5)
The proof is given in Supp. Sec. A.
@@
The theorem above suggests that a safe denoiser can be constructed similarly to CFG. In our case, the denoiser is penalized is determined by $\beta^{*}(\mathbf{x}_{t})$, designed to increase when $\mathbf{x}_{t}$ is likely unsafe. Specifically, a term in the numerator, $\int p_{\text{unsafe}}(\mathbf{x})q_{t}(\mathbf{x}_{t}|\mathbf{x})\mathop{}\!\mathrm{d}\mathbf{x}$, grows as the likelihood of $\mathbf{x}_{t}$ being unsafe increases. In contrast, the denominator grows as the likelihood of $\mathbf{x}_{t}$ being safe increases. Consequently, $\beta^{*}(\mathbf{x}_{t})$ decreases as $\mathbf{x}_{t}$ becomes more likely to be safe. This indicates that our theoretically derived $\beta^{*}(\mathbf{x}_{t})$ shares a similar intuition to the adaptive weight $\mu$ observed in SLD, but correctly aligns with the intended penalty mechanism. In other words, if $\mathbf{x}_{t}$ is more unsafe than $\mathbf{\tilde{x}}_{t}$, then the trajectory of $\mathbf{x}_{t}$ is more penalized than that of $\mathbf{\tilde{x}}_{t}$, i.e., $\beta^{*}(\mathbf{x}_{t})>\beta^{*}(\mathbf{\tilde{x}}_{t})$.
@@
To provide more intuition on the role of the weight in our theorem, we vary the values that the weight can take and show the corresponding samples. In Figure 3-(a), we observe that when safety is considered less rigorously than the measure of $\beta{{}^{*}}(\mathbf{x}_{t})$, some samples reside within the unsafe region. In contrast, Figure 3-(b) demonstrates that by doubling the safety threshold, both the unsafe region and its immediate surroundings are effectively avoided. However, in Figure 3-(c), we observe that the samples from our safe denoiser do not cover the entire safe regions in the data distribution.

### v4

@@ Abstract
There is growing concern over the safety of powerful diffusion models, as they are often misused to produce inappropriate, not-safe-for-work content or generate copyrighted material or data of individuals who wish to be forgotten. Many existing methods tackle these issues by heavily relying on text-based negative prompts or retraining the model to eliminate certain features or samples. In this paper, we take a radically different approach, directly modifying the sampling trajectory by leveraging a negation set (e.g., unsafe images, copyrighted data, or private data) to avoid specific regions of data distribution, without needing to retrain or fine-tune the model. We formally derive the relationship between the expected denoised samples that are safe and those that are unsafe, leading to our safe denoiser, which ensures its final samples are away from the area to be negated. We achieve state-of-the-art safety in large-scale datasets such as the CoPro dataset while enabling significantly more cost-effective sampling than existing methodologies.
@@
(a) Nudity (Top) &\& Violence (Bottom)
(b) Data Memorization
Figure 1: Our method Safe Denoiser against existing methods. (a) Our method, incorporated with SAFREE yoon2024safree and SLD schramowski2023safe , does not generate inappropriate images. (b) Our method mitigates the memorization issue by negating the real image, resulting in a novel image with features similar to those in the real image in hair colors or outfits.
@@ 1 Introduction
Diffusion models (DMs) have become leading generative models, excelling in generation tasks like text-to-image rombach2022high , audio kong2021diffwave , video bartal2024lumiere , and protein design watson2023denovo , thanks to their flexible and controllable sampling dhariwal2021diffusion ; ho2021classifier . However, growing concerns over unsafe content — such as not-safe-for-work(NSFW) imagery (Figure 1(a)), copyright violations, and potential misuse — highlight the need for safety. The key challenge is mitigating these risks without compromising model utility or creativity.
Mainstream mitigation strategies for issues like NSFW content or unwanted concept removal rely on text-based guidance gandikota2023erasing ; ban2024understanding ; yoon2024safree or fine-tuning for unlearning gandikota2023erasing ; gandikota2023unified ; gong2024reliable ; kim2024race . Text-based methods require iterative, expert-crafted negative prompts (schramowski2023safe, ), which may not generalize well, while fine-tuning is resource-intensive and risks catastrophic forgetting or degraded performance on desired tasks.
Other significant safety concerns involve the DMs’ capacity to reproduce copyrighted content and their generation of data pertaining to individuals in Figure 1(b) who wish to be excluded. These issues are often linked to the models’ remarkable ability to memorize training data carlini2023extracting . While techniques like differentially private training dockhorn2023differentially ; liu2024differentially can formally limit memorization by adding noise during the training process, they often result in a noticeable degradation in generation quality, which can be particularly prohibitive for applications demanding high-fidelity outputs.
We propose a safe denoiser (defined in Definition 3.1) that modifies sampling trajectories such that the resulting samples are drawn from a safe distribution (shown in Figure 2). The intuion comes from our Theorem. 3.1, where the safe denoiser steers generation away from unsafe regions, ensuring theoretical safety. We develop a practical algorithm (Algorithm 1) based off of our theorem, which can be used standalone or combined with negative prompting to enhance safety in text-to-image generation. Our method achieves state-of-the-art performance on concept erasing, class removal, and unconditional image generation tasks.
2 Preliminary
@@
DMs generate samples through iterative decoding starting from random noise to data. This iterative process is a reverse of the forward data corruption process, $\mathbf{x}_{t}=\alpha_{t}\mathbf{x}+\sigma_{t}\bm{\epsilon}$, where $\mathbf{x}\sim p_{\text{data}}(\mathbf{x})$ $\bm{\epsilon}\sim\mathcal{N}(0,I)$ which results in a perturbation kernel: $q_{t}(\mathbf{x}_{t}|\mathbf{x})=\mathcal{N}(\mathbf{x}_{t};\alpha_{t}\mathbf{x},\sigma_{t}^{2}I)$. The specific choice of the coefficients $\alpha_{t}$ and $\sigma_{t}$ determines a different variant of DMs: popular examples include Denoising Diffusion Probabilistic Models (DDPM) ho2020denoising , Elucidating Diffusion Models (EDM) karras2022elucidating , or Flow Matching lipman2022flow . Regardless of whether the model is trained with noise-prediction ho2020denoising , data-prediction karras2022elucidating , or velocity-prediction salimans2022progressive ; lipman2022flow , these approaches are fundamentally equivalent kingma2021variational ; kim2021soft . This paper adopts the data-prediction framework due to its most intuitive interpretation. In data-prediction, the model approximates the denoiser function, defined by $\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]:=\int\mathbf{x}\frac{p_{\text{data}}(\mathbf{x})q_{t}(\mathbf{x}_{t}|\mathbf{x})}{p_{\text{data},t}(\mathbf{x}_{t})}\mathop{}\!\mathrm{d}\mathbf{x}\approx\frac{1}{\alpha_{t}}(\mathbf{x}_{t}-\sigma_{t}\bm{\epsilon}_{\bm{\theta}})$, where $p_{\text{data},t}(\mathbf{x}_{t})$ is a marginal distribution of diffusion process at $t$, and $\bm{\epsilon}_{\bm{\theta}}$ is the noise-prediction.
@@
DMs can be guided to produce samples dhariwal2021diffusion ; kim2022refining that adhere more closely to a desired condition denoted by $\mathbf{c}$. A common approach in modern DMs is classifier-free guidance (CFG) ho2021classifier . The model is trained to learn both the unconditional denoiser $\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]$ and the conitional denoiser $\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t},\mathbf{c}]$. The CFG modifies the sampling trajectory by
$\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]+\lambda\big(\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t},\mathbf{c}]-\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]\big)$
@@
allowing stronger alignment of the sample with the prompt $\mathbf{c}$ via the scale $\lambda$. The purpose of the additional term is to guide the trajectory in the sharpening direction toward a desired condition $\mathbf{c}$.
When there are unsafe words in the input text prompt, SAFREE (yoon2024safree, ) detects unsafe words (tokens) and modifies the unsafe token embeddings. It filters out undesirable concepts with
$\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]+\underbrace{\lambda(\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t},\tilde{\mathbf{c}}_{+}]-\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}])}_{\text{SAFREE}},$ (1)
where $\tilde{\mathbf{c}}_{+}$ is a modified prompt embeddings. This altered prompt embedding steers the generation process away from the predefined unsafe concepts.
@@
Another way of negating unsafe concepts is using negative guidance liu2022compositional . It reverses the CFG gradient direction for an undesired prompt denoted by $\mathbf{c}_{-}$. Formally, one replaces the standard CFG with
@@
where $\mathbf{c}_{+}$ denotes a positive condition and $\mathbf{c}_{-}$ represents a negative context that we want to avoid.
On the line of negative prompting, Safe Latent Diffusion (SLD) schramowski2023safe introduces a guidance by
$\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]+\underbrace{\lambda(\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t},\mathbf{c}_{+}]-\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}])}_{\text{CFG}}-\underbrace{\mu(\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t},\tilde{\mathbf{c}}_{-}]-\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}])}_{\text{SLD}},$ (2)
[…]
3 Method
@@
Figure 3: Effect of the weight value in Theorem. 3.1. (a) If we use half the theoretical weight value, samples generated by our weak safe denoiser also cover the unsafe region (i.e., red dots appearing in the blue area). (b) When we use the theoretical value, the samples avoid unsafe regions while covering the whole safe area. (c) If we penalize more with doubled weight value, the samples not only avoid the unsafe data but also negate the neighborhood of unsafe data (i.e., there are no red dots in the black area).
Text-based prompts (like $\mathbf{c}_{-}$ or $\tilde{\mathbf{c}}_{-}$) rely on limited, user-selected words and may miss undesired content. To address this, we introduce a method that offers safety guarantees on generated images, which can be combined with existing text-based safety approaches.
3.1 Safe Denoiser
@@
We first define an indicator function, $1_{\text{safe}}(\mathbf{x})$, taking the value of $1$ if $\mathbf{x}$ is safe and 0 if not. Similarly, we define an indicator function, $1_{\text{unsafe}}(\mathbf{x})$ taking the value of $1$ if $\mathbf{x}$ is unsafe and 0 if not. These indicator functions are the partition of the unity, resulting in $1=1_{\text{safe}}(\mathbf{x})+1_{\text{unsafe}}(\mathbf{x})$ for all $\mathbf{x}\in\text{supp}(p_{\text{data}})$. Then, we define the following concepts.
@@
Definition 3.1. The unnormalized density of the safe distribution $p_{\text{safe}}(\mathbf{x})$ is $1_{\text{safe}}(\mathbf{x})p_{\text{data}}(\mathbf{x})$. The safe denoiser is defined by $\mathbb{E}_{\text{safe}}[\mathbf{x}|\mathbf{x}_{t}]=\int\mathbf{x}\frac{p_{\text{safe}}(\mathbf{x})q_{t}(\mathbf{x}_{t}|\mathbf{x})}{p_{\text{safe},t}(\mathbf{x}_{t})}\mathop{}\!\mathrm{d}\mathbf{x},$ where $p_{\text{safe},t}(\mathbf{x}_{t})$ is the marginal distribution of the diffusion process (at time $t$) starting from the safe distribution. Analogously, the unnormalized density of the unsafe distribution $p_{\text{unsafe}}(\mathbf{x})$ is $1_{\text{unsafe}}(\mathbf{x})p_{\text{data}}(\mathbf{x})$. The unsafe denoiser is $\mathbb{E}_{\text{unsafe}}[\mathbf{x}|\mathbf{x}_{t}]=\int\mathbf{x}\frac{p_{\text{unsafe}}(\mathbf{x})q_{t}(\mathbf{x}_{t}|\mathbf{x})}{p_{\text{unsafe},t}(\mathbf{x}_{t})}\mathop{}\!\mathrm{d}\mathbf{x},$ (3) where $p_{\text{unsafe},t}(\mathbf{x}_{t})$ is the marginal distribution of the diffusion process (at $t$) starting from the unsafe distribution.
@@
Our interest is to obtain $\mathbb{E}_{\text{safe}}[\mathbf{x}|\mathbf{x}_{t}]$ given the data denoiser $\mathbb{E}_{\text{data}}[\mathbf{x}|\mathbf{x}_{t}]$. The theorem below describes the relationship between our safe denoiser and the data denoiser. The proof is given in Appendix A.
Theorem 3.2. Suppose that $\mathbb{E}_{\textup{data}}[\mathbf{x}|\mathbf{x}_{t}]$, $\mathbb{E}_{\textup{safe}}[\mathbf{x}|\mathbf{x}_{t}]$, and $\mathbb{E}_{\textup{unsafe}}[\mathbf{x}|\mathbf{x}_{t}]$ are the data denoiser, the safe denoiser, and the unsafe denoiser. Then, $\mathbb{E}_{\textup{safe}}[\mathbf{x}|\mathbf{x}_{t}]=\mathbb{E}_{\textup{data}}[\mathbf{x}|\mathbf{x}_{t}]+\beta^{*}(\mathbf{x}_{t})\big(\mathbb{E}_{\textup{data}}[\mathbf{x}|\mathbf{x}_{t}]-\mathbb{E}_{\textup{unsafe}}[\mathbf{x}|\mathbf{x}_{t}]\big)$ (4) for a weight defined by $\beta^{*}(\mathbf{x}_{t})=\frac{Z_{\textup{unsafe}}p_{\textup{unsafe},t}(\mathbf{x}_{t})}{Z_{\textup{safe}}p_{\textup{safe},t}(\mathbf{x}_{t})},$ (5) where $Z_{\textup{safe}}:=\int 1_{\textup{safe}}(\mathbf{x})p_{\textup{data}}(\mathbf{x})\mathop{}\!\mathrm{d}\mathbf{x}$ and $Z_{\textup{unsafe}}:=\int 1_{\textup{unsafe}}(\mathbf{x})p_{\textup{data}}(\mathbf{x})\mathop{}\!\mathrm{d}\mathbf{x}$ are normalizing constants of unnormalized densities of safe and unsafe distributions, respectively.
@@
The theorem above suggests that a safe denoiser can be constructed similarly to CFG. In our case, the denoiser is penalized by $\beta^{*}(\mathbf{x}_{t})$, designed to increase when $\mathbf{x}_{t}$ is likely unsafe. Specifically, a term in the numerator, $p_{\text{unsafe},t}(\mathbf{x}_{t})=\int p_{\text{unsafe}}(\mathbf{x})q_{t}(\mathbf{x}_{t}|\mathbf{x})\mathop{}\!\mathrm{d}\mathbf{x}$, grows as the likelihood of $\mathbf{x}_{t}$ being unsafe increases. In contrast, the denominator grows as the likelihood of $\mathbf{x}_{t}$ being safe increases. Consequently, $\beta^{*}(\mathbf{x}_{t})$ decreases as $\mathbf{x}_{t}$ becomes more likely to be safe. This indicates that our $\beta^{*}(\mathbf{x}_{t})$ shares a similar intuition to the adaptive weight $\mu$ observed in SLD, but correctly aligns with the intended penalty mechanism. In other words, if $\mathbf{x}_{t}$ is more unsafe than $\mathbf{\tilde{x}}_{t}$, then the trajectory of $\mathbf{x}_{t}$ is more penalized than that of $\mathbf{\tilde{x}}_{t}$, i.e., $\beta^{*}(\mathbf{x}_{t})>\beta^{*}(\mathbf{\tilde{x}}_{t})$.
@@
To provide more intuition on the role of the weight in our theorem, we vary the values that the weight can take and show the corresponding samples. In Figure 3(a), we observe that when safety is considered less rigorously than the measure of $\beta{{}^{*}}(\mathbf{x}_{t})$, some samples reside within the unsafe region. In contrast, Figure 3(b) demonstrates that by doubling the safety threshold, both the unsafe region and its immediate surroundings are effectively avoided. However, in Figure 3(c), we observe that the samples from our safe denoiser do not cover the entire safe regions in the data distribution.